OpenStudy (anonymous):

Evaluate the limit lim as a approaches 49 (49-a)/7-radical(a)

OpenStudy (anonymous):



OpenStudy (loser66):

multiple both numerator and denominator by \(7+\sqrt a\)

OpenStudy (anonymous):

consider \(49-a=(7+\sqrt{a})(7-\sqrt{a})\)

OpenStudy (anonymous):

that makes sense @oldrin.bataku , so then would i just cancel out the 7- radical(a)

OpenStudy (loser66):

yes, he is smarter than me. hehehe... his way is good

OpenStudy (anonymous):

well your smart too @Loser66

OpenStudy (xapproachesinfinity):

same thing! you can go either way

OpenStudy (anonymous):

nevermind i got it thnx @oldrin.bataku

OpenStudy (xapproachesinfinity):

yeah i know just commenting on what @Loser66 said